# A look on equations describing pseudospherical surfaces

### v1

@@ 2 Equations describing pseudospherical surfaces
Remark 2.3.
The theory of PSS equations was conceived only for equations, or even systems of equations, in mind. Additional information, such as boundary or initial conditions, was not considered for a long time.
Remark 2.4.
@@
The theory of PSS equations has traditionally been formulated under the assumption that the solutions defining the first fundamental form possess $C^{\infty}$ regularity. Although this smoothness requirement is not always explicitly stated, it can be traced back to the foundational work of Chern and Tenenblat [8, page 55, second paragraph]. In contrast, the discussion in the preceding paragraph highlights a tension between this framework and the class of solutions considered in [10, 11, 13], where the smoothness assumption is incompatible with the finite regularity typically associated with Cauchy problems. Consequently, surfaces potentially generated by solutions exhibiting wave breaking remained unexplored within this geometric context until [20].
@@ 3 Finite regularity of the metric
In this section we address problems concerning PSS equations and some developments made by the author in [16, 19, 20, 21, 22, 37]. The CH equation was the main motivation for such improvements and thus, parts of the text are now more focused on its literature. For a review on some aspects of the CH equation, see [18].
@@
This distinction motivates a deeper investigation into the relationship between regular curves and the graphs of solutions of the CH equation. More precisely, one may ask under which conditions a regular curve in space can uniquely determine such a graph, together with a compatible set of one-forms. In certain cases, uniqueness can indeed be established. This situation essentially corresponds to solving the CH equation with prescribed initial data $u_{0}$, which naturally determines a curve of the form $(x,0,u_{0}(x))$. However, the requirement of uniqueness imposes constraints on the class of curves that can be considered.
@@
As previously noted, solutions to Cauchy problems for the CH equation do not always exhibit smoothness [10, 11, 13], that is, their solutions may not be $C^{\infty}$. This raises the issue of what minimal regularity is necessary to ensure the well-definition of the associated one-forms.
@@
Although the theory of PSS equations generally assumes smooth solutions, this requirement is often implicit. Exceptions are rare, with the work [37] standing out as a notable example. There, PSS surfaces derived from solutions to a specific Cauchy problem were considered, but the solutions in question were analytic ($C^{\omega}$), and hence smooth. Nevertheless, [37] also demonstrated the existence of a strip in which the one-forms are defined and the condition $\omega_{1}\wedge\omega_{2}\neq 0$ is satisfied.
@@
It is worth recalling that, for a solution to define a first fundamental form of a surface, the associated one-forms must be linearly independent over simply connected domains in the plane [9, Theorem 4.39]. This requirement prompts further examination of how the geometry of the solution’s graph evolves, particularly with respect to the domain where a PSS structure can be meaningfully defined.
@@
Solutions of the CH equation may develop singularities in finite time. For example, Constantin and Escher [11, Example 4.3] demonstrated that even smooth initial data may lead to wave breaking in finite time. The formation of such singularities is typically expressed through the blow-up of the slope of the solution.
@@
The source of the problems above is precisely one: the space where a given solution is considered. From a geometric perspective, provided that the one-forms (2.0.2) are $C^{1}$ (see [23, Theorem 10-19, page 232] and also [23, Theorem 10-18, page 232]), we do have an abstract surface defined. Therefore, as long as we have a solution belonging to a function space respecting this constraint, the theory developed so far continues working.
With all of this in mind, we can look for minor modifications in the notion of PSS equations and generic solutions, e.g., see [20, 21].
From now on, we closely follows the ideas presented in [20]. Our presentation is more concerned with the geometric nature of the results and we avoid technicalities with functional analysis. However, we guide the reader for [20] to clarify certain aspects on the function spaces involved, such as the Sobolev spaces $H^{s}(\mathbb{R})$ and other related tools.
Definition 3.1.
@@
  • the forms (2.0.2) satisfy the structure equations (2.0.3) of a pseudospherical surface of Gaussian curvature ${\cal K}=-1$;
@@
If the function space is the space of $C^{\infty}$ functions, or it is not exactly or clearly defined, and no confusion is possible, we then say PSS equation instead of ${\cal B}-$PSS equation.
@@
The one-forms involved in Definition 2.1 are pull-backs of differential forms on jet spaces; see [30, Section II].
Definition 3.2.
(Generic solution) A function $u:U\rightarrow\mathbb{R}$ is called a generic solution for the ${\cal B}-$PSS equation (2.0.1) provided that:
  1. a)
@@
  3. c)
    the one-forms (2.0.2) are $C^{k}$ on $U$;
  4. d)
    there exists at least a simply connected open set $\Omega\subseteq U$ such that $\omega_{1}\wedge\omega_{2}\big{|}_{p}\neq 0$, for each $p\in\Omega$.
@@
Additionally, all forms we will deal with herein have to be understood in a similar fashion to the above comment. For a better discussion, see [31, pages 77–78].
@@
Assume that $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$ is a non-trivial solution of the CH equation (1.0.6). If $u_{x}\neq 0$ on some open set $\Omega$, then we cannot have $m=c\neq 0$, where $c$ is a constant.
Proof.
Suppose $m=c\neq 0$ is a constant on some open and non-empty set $\Omega$, then necessarily $m\neq 0$ and (1.0.6) implies (2.0.11). On the other hand, if $u$ is a solution of (1.0.6) such that $u_{x}=0$ on an open set $\Omega$, then $u_{x}=u_{xx}=u_{xxx}=u_{txx}=0$ and $u_{t}=0$ on $\Omega$, which, once substituted into (1.0.6), implies $u(x,t)=c$. ∎
Theorem 3.1.
@@
Proof.
Let $u_{0}\in H^{4}(\mathbb{R})$. From [10, Proposition 2.7] with $s=4$ we can find a number $T>0$, uniquely determined by $u_{0}$, and determine a unique function $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$ that is a solution of (1.0.5). The lifespan $T$ determined by the initial datum uniquely defines the strip mentioned in the statement of the theorem, whereas the solution $u$, granted by [10, Proposition 2.7], ensures that (2.0.9) are $C^{1}$ one-forms defined on ${\cal S}$.
Since $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$, we have $u_{x}\in H^{3}(\mathbb{R})$ for each fixed $t\in[0,T)$. Moreover, since the $H^{1}(\mathbb{R})$-norm of the solutions of the CH equation is invariant, and the fact that $u_{0}$ is non-trivial, imply that for any value of $t$ for which the solution exists, we have
$\|u(\cdot,t)\|_{H^{1}(\mathbb{R})}=\|u_{0}(\cdot)\|_{H^{1}(\mathbb{R})}>0,$ (3.0.1)
@@
$\lim_{|x|\rightarrow\infty}u(x,t)=0\quad\text{and}\quad\lim_{|x|\rightarrow\infty}u_{x}(x,t)=0.$ (3.0.2)
@@
By Lemma 3.1, $m$ cannot be constant on $B_{\epsilon}(b_{t},t)$. As a result, for some $(y_{0},t_{0})\in B_{\epsilon}(b_{t},t)$ we have $m(y_{0},t_{0})\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$. Then, by continuity, we can find $\epsilon^{\prime}\in(0,\epsilon]$ such that $m(x,t)\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$, for all $(x,t)\in B_{\epsilon^{\prime}}(y_{0},t_{0})\subseteq B_{\epsilon}(b_{t},t)$. Let $B_{1}:=B_{\epsilon^{\prime}}(y_{0},t_{0})$. The same argument shows the existence of another disc $B_{2}\subseteq B_{\epsilon}(a_{t},t)$ where $m\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$.
Therefore, by the first part of the theorem, the one-forms (2.0.8) are defined on these discs, while the conditions above say that $\omega_{1}\wedge\omega_{2}\big{|}_{B_{1}}\neq 0$, $\omega_{1}\wedge\omega_{2}\big{|}_{B_{2}}\neq 0$, and they do not vanish at any point of $B_{1}\cup B_{2}$.
Let us now prove the claim by showing that we cannot have $h(t)\geq 0$. If we had $h(t_{0})\geq 0$ for some $t_{0}\in[0,T)$, then either $u_{x}(x,t_{0})\geq 0$ or $u_{x}(x,t_{0})\leq 0$ for all $x\in\mathbb{R}$, meaning that $u$ is either non-decreasing or non-increasing. Therefore, (3.0.2) implies $u(\cdot,t_{0})\equiv 0$, which contradicts (3.0.1). ∎
We have just shown that any non-trivial initial datum defines uniquely a strip ${\cal S}$ containing open sets that can be endowed with the structure of a PSS. Our task now is to look for singularities of the metric defined by the forms (2.0.8). By a singularity, we mean a point $p$ for which the metric is singular, in the sense that it is no longer a positive definite bilinear form. This situation occurs whenever the one-forms $\omega_{1}$ and $\omega_{2}$ are not linearly independent at $p$, that is, $\omega_{1}\wedge\omega_{2}\big{|}_{p}=0$.
Theorem 3.2.
Assume that $u_{0}\in H^{4}(\mathbb{R})$ is a non-trivial initial datum, $u$ the corresponding solution of (1.0.5), and let ${\cal S}$ be the strip determined in Theorem 3.1.
  1. a)
@@
Proof.
By Example 2.2, $u$ is a generic solution of the CH equation provided that $u_{x}=0$ or $m=const.$ For convenience, let us consider the CH equation in the form (1.0.6). If we assume that (2.0.10) holds on some open and non-empty set $\Omega$, then necessarily $m\neq 0$ and (1.0.6) implies (2.0.11). On the other hand, Lemma 3.1 implies that if $m$ is constant on some open set, then $u(x,t)$ coincides with $m$ within the same open set.
Let us show the existence of at least one point $c_{t}$ such that $u_{x}(c_{t},t)=0$ for each $t\in[0,T)$. Let $h$, $a_{t}$ and $b_{t}$ be as in the proof of Theorem 3.1. The two points $a_{t}$ and $b_{t}$ cannot be the same and the intermediate value theorem ensures the existence of $c_{t}\in(a_{t},b_{t})$ such that $u_{x}(c_{t},t)=0$.
Finally, let us prove that $\Omega$ must be a proper subset of ${\cal S}$. In fact, if we had $\Omega={\cal S}$, then we would be forced to conclude that $c=0$ from (3.0.2), which would then lead to a contradiction with (3.0.1). ∎
4 Final remarks

### v4

@@ 2 Equations describing pseudospherical surfaces
It is also worth mentioning that even though geometric integrability and AKNS integrability are related and shared by a number of equations, they are not equivalent. The Cavalcanti and Tenenblat equation
$u_{t}=(u_{x}^{-1/2})_{xx}+u_{x}^{3/2},$
discovered by Cavalcanti and Tenenblat [4], is the compatibility condition for the one-forms
$\begin{array}[]{lcl}\omega_{1}&=&\eta\sinh{u}\,dx+\eta\Big((u_{x}^{-1/2})_{x}\cosh{u}+(u_{x}^{1/2}-\eta u_{x}^{-1/2})\sinh{u}\Big)dt,\\ \\ \omega_{2}&=&\eta dx-\eta^{2}u_{x}^{-1/2}dt,\\ \\ \omega_{3}&=&\eta\cosh{u}\,dx+\eta\Big((u_{x}^{-1/2})_{x}\sinh{u}+(u_{x}^{1/2}-\eta u_{x}^{-1/2})\cosh{u}\Big)dt,\end{array}$
satisfies the structure equations (11), but it is not of the AKNS type [32, page 79].
Remark 2.7 (An equation describing pseudospherical surfaces is not necessarily geometrically integrable).
If an equation is geometrically integrable, then it describes one-parameter families of PSS. This is due to the presence of a (at least one) parameter in the one-forms that cannot be removed by a gauge transformation $X\mapsto SXS^{-1}$, $T\mapsto STS^{-1}$, where $X$ and $T$ satisfy the ZCR (8). We observe that such a gauge transformation would change the corresponding one-forms but does not affect geometric integrability.
Let us consider the one-forms (see [39, Example 2.8])
$\begin{array}[]{lcl}\omega_{1}&=&(u-u_{xx})dx+(u_{x}^{2}-2uu_{x}+uu_{xx})dt,\\ \\ \omega_{2}&=&\Big(\mu(u-u_{xx})\pm 2\sqrt{1+\mu^{2}}\Big)dx+\mu(u_{x}^{2}-2uu_{x}+uu_{xx})dt,\\ \\ \omega_{3}&=&\Big(\pm\sqrt{1+\mu^{2}}(u-u_{xx})+2\mu\Big)dx\pm\sqrt{1+\mu^{2}}(u_{x}^{2}-2uu_{x}+uu_{xx})dt.\end{array}$ (22)
A straightforward calculation shows that
$\begin{array}[]{lcl}d\omega_{1}-\omega_{3}\wedge\omega_{2}&=&{\Big(u_{t}-u_{txx}+4uu_{x}-3u_{x}u_{xx}-uu_{xxx}\Big)dx\wedge dt,}\\ \\ d\omega_{2}-\omega_{1}\wedge\omega_{3}&=&0,\\ \\ d\omega_{3}-\omega_{1}\wedge\omega_{2}&=&{-\Big(u_{t}-u_{txx}+4uu_{x}-3u_{x}u_{xx}-uu_{xxx}\Big)dx\wedge dt}.\end{array}$
Therefore, the one-forms (22) satisfy the structure equations for a PSS provided that $u$ is a solution of the Degasperis-Procesi equation
$u_{t}-u_{txx}+4uu_{x}=3u_{x}u_{xx}+uu_{xxx}.$
The presence of the parameter $\mu$ in (22) might suggest that the Degasperis-Procesi equation is geometrically integrable. However, this is not the case, at least with the forms (22).
Under a gauge transformation, see [18, Section 8], the one-forms (22) are equivalent to
$\begin{array}[]{lcl}\theta_{1}&=&-2dx,\\ \\ \theta_{2}&=&{\Big(1+\frac{u-u_{xx}}{2}\Big)dx+\frac{u_{x}^{2}-2uu_{x}+uu_{xx}}{2}dt},\\ \\ \theta_{3}&=&{\Big(1+\frac{u-u_{xx}}{2}\Big)dx+\frac{u_{x}^{2}-2uu_{x}+uu_{xx}}{2}dt,}\end{array}$
which does not depend on any parameter. It is important to mention that this fact does not imply that the Degasperis-Procesi equation is not geometrically integrable! It does imply that it is not geometrically integrable, considering (22), because the parameter in those forms can be removed. A discussion about geometric integrability of the Degasperis-Procesi equation can be found in [18, Discussion].
Remark 2.8 (PSS equations are a proper category).
The set of PSS equations is larger than the set of integrable equations à la AKNS integrability. In fact, those PSS equations that are integrable (in any sense we may consider) represent a small, but quite important, number of what today we know to be PSS equations. This distinction makes evident that the existence of equations being the compatibility condition for the structural equations (11) reveals a universal geometric feature, whereas the integrability aspect, despite its relevance, is a specific subcase.
Remark 2.9 (PSS from smooth solutions).
None of the definitions 2.1 or 2.2 made reference of the regularity of the objects involved. This is intentional: they follow the spirit of papers in geometry, where usually objects are assumed to be $C^{\infty}$. In particular, all solutions involved are $C^{\infty}$ and so are the forms (10).
Remark 2.10 (PSS equations and Cauchy problems).
The developments in the field of PSS equations had been restricted to equations or systems of equations. Additional information, such as boundary or initial conditions, was not considered until a few years ago. This point will be better explored in the next section.
Remark 2.11 (Limitations of the theory of PSS equations).
@@
The theory of PSS equations has traditionally been formulated under the assumption that the solutions defining the first fundamental form possess $C^{\infty}$ regularity. Although this smoothness requirement is not always explicitly stated, it can be traced back to the foundational work of Chern and Tenenblat [5, page 55, second paragraph]. In contrast, the discussion in the preceding remark highlights a tension between this framework and the class of solutions considered in [8, 10, 7], where the smoothness assumption is incompatible with the finite regularity typically associated with Cauchy problems. Consequently, surfaces potentially generated by solutions exhibiting wave breaking remained unexplored within this geometric context until [16].
@@ 3 Finite regularity of the metric
In this section, we address problems concerning PSS equations and some developments made by the author in [29, 19, 16, 18, 17, 14]. The CH equation was the main motivation for such improvements, and thus, parts of the text are now more focused on its literature. For a review on some aspects of the CH equation, see [15].
@@
This distinction motivates a deeper investigation into the relationship between certain regular curves and the graphs of solutions of the CH equation. More precisely, one may ask under which conditions a regular curve $\gamma(x)=(x,0,u_{0}(x))$ in space can uniquely determine such a graph, together with a compatible set of one-forms. In certain cases, uniqueness can indeed be established. This situation essentially corresponds to solving the CH equation with prescribed initial data $u_{0}$, which naturally determines a curve of the form $(x,0,u_{0}(x))$. However, the requirement of uniqueness imposes constraints on the class of curves of this type that can be considered.
@@
As previously noted, solutions to Cauchy problems for the CH equation do not always exhibit smoothness [8, 10, 7], that is, their solutions may not be $C^{\infty}$. Some issues naturally arise from this observation:
  • are definitions 2.2 and 2.1 still valid in the absence of $C^{\infty}$ regularity?
  • if the answer is positive, what is the minimal regularity necessary to ensure the well-definition of the associated one-forms?
@@
In light of the preceding comments, from a geometric perspective, PSS equations are not usually dealt with other conditions or restrictions, such as initial conditions. Exceptions are rare, and to the best of the author’s knowledge, [14] stands out as a first example of the Cauchy problem being considered in conjunction with the geometry of PSS equations. There, PSS surfaces derived from solutions to a specific Cauchy problem were considered, but the solutions in question were analytic ($C^{\omega}$), and hence smooth. Nevertheless, [14] also demonstrated the existence of a strip in which the one-forms are defined and the condition $\omega_{1}\wedge\omega_{2}\neq 0$ is satisfied.
@@
It is worth recalling that, for a solution to define a first fundamental form of a surface, the associated one-forms must be linearly independent over simply connected domains in the plane [6, Theorem 4.39]. This requirement prompts further examination of how the geometry of the solution’s graph evolves, particularly with respect to the domain where a PSS structure can be meaningfully defined.
Despite the considerations in [14], the solutions considered therein were still $C^{\infty}$, so that the framework proposed by Tenenblat and co-workers over the years was still valid. The situation, however, is significantly different when solutions of the CH equation are taken into consideration.
@@
Cauchy problems involving the CH equation are usually established with finite regularity, since they are $C^{1}$ in time [8, 10, 7, 9]. Moreover, solutions of the CH equation may develop singularities in finite time. For example, Constantin and Escher [10, Example 4.3] demonstrated that even smooth initial data may lead to wave breaking in finite time. The formation of such singularities is typically expressed through the blow-up of the slope of the solution. The presence of finite regularity certainly brings some difficulties to Definition 2.1, but with Definition 2.2 the issue becomes dramatic, since it is not clear if we still have an intrinsic geometry defined by (10)–(11) whenever the solution of the equation has finite regularity.
@@
The initial data (which geometrically gives a curve) provides the regularity of the solution because it determines the space function where the solutions belong to. This is precisely the source of the problems mentioned above: the space where a given solution is considered. From a geometric perspective, provided that the one-forms (10) are $C^{1}$ (see [21, Theorem 10-19, page 232] and also [21, Theorem 10-18, page 232]), we do have an abstract surface defined. Therefore, as long as we have a solution belonging to a function space respecting this constraint, the theory developed so far continues to work.
With all of this in mind, we have a clue to modify Definition 2.1 and make it compatible when we are out of $C^{\infty}$ structures: the definition should take into account where the solution is. This is, in fact, done in its original assumption, since it implicitly assumes $C^{\infty}$ solutions. If we allow other solutions, we can then bring modifications in the notion of PSS equations and generic solutions, e.g., see [16, 18].
[…]
@@
  • the forms (10) satisfy the structure equations (11) of a pseudospherical surface of Gaussian curvature ${\cal K}=-1$;
@@
[…]
@@
The one-forms involved in Definition 3.1 are pullbacks of differential forms on jet spaces; see [31, Section II].
Remark 3.3.
One might wonder whether the order of the derivatives appearing in the functions $f_{ij}$ in Definition 3.1 is related to the order of the equation. Most commonly, they are of a smaller order. However, this is not a rule. For example, let us consider the one-forms
$\begin{array}[]{lcl}\omega_{1}&=&-2dx,\\ \\ \omega_{2}=\omega_{3}&=&{\Big(1-2u+u_{x}+2u_{xx}-u_{xxx}\Big)dx}\\ \\ &+&{\Big(16u_{x}^{2}-16uu_{x}+16uu_{xx}-16u_{x}u_{xx}-4uu_{xxx}+2u_{xx}^{2}+2u_{x}u_{xxx}\Big)dt},\end{array}$
which were reported in [11, Theorem 2.1].
Their coefficients depend on up to third-order derivatives, and a lengthy calculation shows that
$\begin{array}[]{lcl}d\omega_{1}&-&\omega_{3}\wedge\omega_{2}\equiv 0,\\ d\omega_{2}&-&\omega_{1}\wedge\omega_{3}=d\omega_{3}-\omega_{1}\wedge\omega_{2}\\ &=&(2-\partial_{x})\Big(u_{t}-u_{txx}-\big(16uu_{x}-8u_{x}u_{xx}+2u_{xx}^{2}-4uu_{xxx}+2u_{x}u_{xxx}\big)\Big)dx\wedge dt.\end{array}$
Therefore, on the solutions of the equation
$u_{t}-u_{txx}=16uu_{x}-8u_{x}u_{xx}+2u_{xx}^{2}-4uu_{xxx}+2u_{x}u_{xxx}$
the one-forms satisfy the structure equations for a PSS.
From now on, we closely follow the ideas presented in [16]. Our presentation is more concerned with the geometric nature of the results and we avoid technicalities with functional analysis. However, we guide the reader for [16] to clarify certain aspects on the function spaces involved, such as the Sobolev spaces $H^{s}(\mathbb{R})$ and other related tools.
Definition 3.4 (Generic solution, [16, Definition 2.2]).
[…]
  1. 1.
@@
  3. 3.
    the one-forms (10) are $C^{k}$ on $U$;
  4. 4.
    there exists at least one simply connected open set $\Omega\subseteq U$ such that $\omega_{1}\wedge\omega_{2}\big|_{p}\neq 0$, for each $p\in\Omega$.
@@
Henceforth, we consider the results in Example 2.4
@@
Assume that $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$ is a non-trivial solution of the CH equation (19). If $u_{x}\neq 0$ on some open set $\Omega$, then we cannot have $m=c\neq 0$, where $c$ is a constant.
Proof 3.7.
Suppose $m=c\neq 0$ is a constant on some open and non-empty set $\Omega$, then necessarily $m\neq 0$ and (19) implies (21). On the other hand, if $u$ is a solution of (19) such that $u_{x}=0$ on an open set $\Omega$, then $u_{x}=u_{xx}=u_{xxx}=u_{txx}=0$ and $u_{t}=0$ on $\Omega$, which, once substituted into (19), implies $u(x,t)=c$.
Theorem 3.8 ([16, Theorem 2.2]).
@@
Proof 3.9.
Let $u_{0}\in H^{4}(\mathbb{R})$. From [8, Proposition 2.7] with $s=4$ we can find a number $T>0$, uniquely determined by $u_{0}$, and determine a unique function
$u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$
that is a solution of (18). The lifespan $T$ determined by the initial datum uniquely defines the strip mentioned in the statement of the theorem, whereas the solution $u$, granted by [8, Proposition 2.7], ensures that (17) are $C^{1}$ one-forms defined on $\cal S$}.
Since $u\in C^{0}(H^{4}(\mathbb{R}),[0,T))\cap C^{1}(H^{3}(\mathbb{R}),[0,T))$, we have $u_{x}\in H^{3}(\mathbb{R})$ for each fixed $t$ in $[0,T)$. Moreover, since the $H^{1}(\mathbb{R})$-norm of the solutions of the CH equation is invariant, and the fact that $u_{0}$ is non-trivial, implies that for any value of $t$ for which the solution exists, we have
$\|u(\cdot,t)\|_{H^{1}(\mathbb{R})}=\|u_{0}(\cdot)\|_{H^{1}(\mathbb{R})}>0,$ (23)
@@
$\lim_{|x|\rightarrow\infty}u(x,t)=0\quad\text{and}\quad\lim_{|x|\rightarrow\infty}u_{x}(x,t)=0.$ (24)
@@
By Lemma 3.6, $m$ cannot be constant on $B_{\epsilon}(b_{t},t)$. As a result, for some $(y_{0},t_{0})$ in $B_{\epsilon}(b_{t},t)$ we have $m(y_{0},t_{0})\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$. Then, by continuity, we can find $\epsilon^{\prime}\in(0,\epsilon]$ such that $m(x,t)\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$, for all $(x,t)\in B_{\epsilon^{\prime}}(y_{0},t_{0})\subseteq B_{\epsilon}(b_{t},t)$. Let $B_{1}:=B_{\epsilon^{\prime}}(y_{0},t_{0})$. The same argument shows the existence of another disc $B_{2}\subseteq B_{\epsilon}(a_{t},t)$ where $m\neq\frac{\lambda}{2}+\frac{1}{2\lambda}$.
Therefore, by the first part of the theorem, the one-forms (16) are defined on these discs, while the conditions above say that $\omega_{1}\wedge\omega_{2}\big|_{B_{1}}\neq 0$, $\omega_{1}\wedge\omega_{2}\big|_{B_{2}}\neq 0$, and they do not vanish at any point of $B_{1}\cup B_{2}$.
Now, we prove the claim by showing that we cannot have $h(t)\geq 0$. If we had $h(t_{0})\geq 0$ for some $t_{0}\in[0,T)$, then either $u_{x}(x,t_{0})\geq 0$ or $u_{x}(x,t_{0})\leq 0$ for all $x\in\mathbb{R}$, meaning that $u$ is either non-decreasing or non-increasing. Therefore, (24) implies $u(\cdot,t_{0})\equiv 0$, which contradicts (23).
We have just shown that any non-trivial initial datum defines uniquely a strip $\cal S$} containing open sets that can be endowed with the structure of a PSS. Our task now is to look for singularities of the metric defined by the forms (16). By a singularity, we mean a point $p$ for which the metric is singular, in the sense that it is no longer a positive definite bilinear form. This situation occurs whenever the one-forms $\omega_{1}$ and $\omega_{2}$ are not linearly independent at $p$, that is, $\omega_{1}\wedge\omega_{2}\big|_{p}=0$.
Theorem 3.10 ([16, Theorem 2.3]).
Assume that $u_{0}\in H^{4}(\mathbb{R})$ is a non-trivial initial datum, $u$ the corresponding solution of (18), and let $\cal S$} be the strip determined in Theorem 3.8.
  1. 1.
@@
Proof 3.11.
From Example 2.4, $u$ is a non-generic solution of the CH equation provided that $u_{x}=0$ or $m=const.$ For convenience, let us consider the CH equation in the form (19). If we assume that (20) holds on some open and non-empty set $\Omega$, then necessarily $m\neq 0$ and (19) implies (21). On the other hand, Lemma 3.6 implies that if $m$ is constant on some open set, then $u(x,t)$ coincides with $m$ within the same open set.
Let us show the existence of at least one point $c_{t}$ such that $u_{x}(c_{t},t)=0$ for each $t\in[0,T)$. Let $h$, $a_{t}$, and $b_{t}$ be as in the proof of Theorem 3.8. The two points $a_{t}$ and $b_{t}$ cannot be the same, and the intermediate value theorem ensures the existence of $c_{t}\in(a_{t},b_{t})$ such that $u_{x}(c_{t},t)=0$.
Finally, let us prove that $\Omega$ must be a proper subset of $\cal S$}. In fact, if we had $\Omega={\cal S}$, then we would be forced to conclude that $c=0$ from (24), which would then lead to a contradiction with (23).
The geometric phenomena described in the Theorems 3.1 and 3.2 are not particular aspects related to the CH equation. They also appear, for instance, in the study of the Degasperis-Procesi equation and its connections with PSS surfaces, see [18].
4 Further remarks